OpenStudy (anonymous):

I need help with this problem please? Find the four real zeros (including multiplicity) of the polynomial f(x)=x^4+3x^3-12x^2-20x+48. Determine the sum of all the zeros of f (x). (Hint: The real zeros are integers that lie between -4 and 4, inclusive.) a. 9 b. -3 c. -5 d. none of these

OpenStudy (ranga):

Are you familiar with rational roots theorem?

OpenStudy (anonymous):

Yes, sort of!

OpenStudy (ranga):

Actually rational roots theorem may not be needed here because they give you a hint that the real zeros are integers that lie between -4 and 4, inclusive. All you have to do is start with x = +1. Put that into x^4+3x^3-12x^2-20x+48 and see if it goes to zero. If it does then x = +1 is a root. If not try x = -1, then x = +2 and x = -2, x = +3, x = -3, x = +4 and x = - 4.

OpenStudy (jdoe0001):

so if x = +1 is a root, then x = 1 => x-1=0 (x-1) = 0 divide your polynomial by x-1, the remainder should be 0, and keep the quotient do the same to find a 2nd root for the quotient

OpenStudy (anonymous):

Okay!

OpenStudy (anonymous):

I think the answer is -3. Is that correct?

OpenStudy (jdoe0001):

hmm dunno.... haven't gotten the roots yet, what roots did you get?

OpenStudy (ranga):

-3 is one of the roots.

OpenStudy (anonymous):

I'm still working on it... I got -3 as a root!

OpenStudy (anonymous):

Thanks ranga!

OpenStudy (ranga):

Since they tell you the roots are all integers in the range -4 to +4, the quickest way to find the roots is to just plug in the 8 integers and see if f(x) = 0.

OpenStudy (anonymous):

Alright! I'll do that!

OpenStudy (ranga):

But you may have to end up doing synthetic division because one of the roots has a multiplicity of 2 I think.

OpenStudy (anonymous):

Okay!

OpenStudy (jdoe0001):

... yeah.... -3 is a root... so \(\bf x^4+3x^3-12x^2-20x+48\implies (x-3)(x^3-12x+16)\)

OpenStudy (anonymous):

Okay, thanks!

OpenStudy (ranga):

Oh, wait a minute. They don't ask you to find all the roots. They give you multiple choices and only one fits the bill.

OpenStudy (anonymous):

Oh, that's right! Sorry I didn't make the answer choices clear!

OpenStudy (ranga):

Then you are done! You found the answer.

OpenStudy (anonymous):

Great! Does that mean that -3 is the answer? =)

OpenStudy (ranga):

yes.

OpenStudy (anonymous):

Alright! =) Both of you were so much help! Thank you!

OpenStudy (ranga):

you are welcome.

OpenStudy (jdoe0001):

hmmm.... I read it as .... find the 4 roots, then sum them up to get an integer for the choices

OpenStudy (ranga):

oops. I should read questions fully before answering. :) Sorry back to the drudgery of finding all roots! :(

OpenStudy (anonymous):

It's okay!

OpenStudy (jdoe0001):

so +4 is another... so \(\bf x^4+3x^3-12x^2-20x+48\\ \quad \\ \implies (x-3)(x+4)(x^2+4x-4)\) and then you'd just have to do the quadratic

OpenStudy (anonymous):

Alright, I'm working it out : )

OpenStudy (ranga):

BTW, if they are just interested in the sum of the roots we can quickly find the sum of the roots without actually finding the roots. For a polynomial, the sum of the roots will be -b/a where a is the leading coefficient and b is the coefficient of the next term. For x^4+3x^3-12x^2-20x+48 a = 1, b = 3 -b/a = -3/1 = -3. So the sum of the roots will be -3. But here they expect you to actually find all 4 roots and add them up.

OpenStudy (anonymous):

Okay, thanks for showing the work =)

OpenStudy (ranga):

you are welcome. Do you know synthetic division? As soon as you found one root, -3, you can do synthetic division, get the quotient, then find the root of the quotient by substituting various x values and as soon as you find another root to do synthetic division, finding the quotient, etc.

OpenStudy (anonymous):

Yes I know synthetic division =)

OpenStudy (ranga):

Then as soon as you found the first root, do synthetic division and find the quotient: x^4+3x^3-12x^2-20x+48 x = -3 is a root. Do synthetic division: | -3 | 1 +3 -12 -20 +48 | -3 0 +36 -48 |________________________________________ 1 0 -12 +16 0 So the quotient is: x^3 - 12x + 16 So x^4+3x^3-12x^2-20x+48 = (x + 3)(x^3 - 12x + 16) To find another root, you can try the few integer values on the quotient. So try x = -4 in x^3 - 12x + 16 and see if it goes to 0.

OpenStudy (anonymous):

Alright, I get that!

OpenStudy (anonymous):

I'm doing x = -4 in x^3 - 12x + 16 now!

OpenStudy (anonymous):

This is a lot of work! :( sorry it's taking so long!

OpenStudy (ranga):

If we put x = -4 in x^3 - 12x + 16 we get (-4)^3 - 12(-4) + 16 = -64 + 48 + 16 = -64 + 64 = 0 Therefore, x = -4 is another root. So do synthetic division of x^3 - 12x + 16 by -4. (note: the x^2 term is missing and so put a zero in its place)

OpenStudy (ranga):

| -4 | 1 0 -12 16 | |_____________________________

OpenStudy (anonymous):

Okay I got (x^3+5x^2+20x+68) remainder 288

OpenStudy (anonymous):

Using synthetic division! ^

OpenStudy (ranga):

way off! Looks like you may have multiplied numbers instead of adding.

OpenStudy (anonymous):

Oh sorry : (

OpenStudy (ranga):

no big deal. give it a go again. drop the 1. multiply 1 and -4 and put it under 0. add. multiply by -4, etc...

OpenStudy (anonymous):

Okay

OpenStudy (ranga):

Do synthetic division of x^3 - 12x + 16 by -4: x^3 - 12x + 16 = x^3 + 0x^2 - 12x + 16 | -4 | 1 0 -12 16 | -4 +16 -16 |_____________________________ 1 -4 4 0 Quotient is: x^2 - 4x + 4

OpenStudy (anonymous):

sorry my internet is crashing! It keeps taking me out of this page

OpenStudy (anonymous):

It's okay, I'll solve the rest, thanks for all the help it is really appreciated =)

OpenStudy (ranga):

x^2 - 4x + 4 can be factored. x^2 - 4x + 4 = x^2 - 2x - 2x + 4 x(x - 2) - 2(x - 2) (x - 2)(x - 2) = (x - 2)^2 So x = 2 is another root and it has a multiplicity of 2 because of (x - 2)^2. That means the root x = 2 occurs twice. So the original f(x) = x^4+3x^3-12x^2-20x+48 = (x - 2)^2 * (x + 3) * (x + 4) The four roots are: -4, -3, +2, +2 Add them up and you get: - 4 + (-3) + 2 + (-2) = -4 - 3 + 2 + 2 = -7 + 4 = -3 As we expected it to be. So the answer is b.

OpenStudy (ranga):

no problem. And you are welcome.